OpenStudy (anonymous):

MEDAL!! Determine the period, domain, range, zeros, and asymptotes (if any). y=-3tan(1/2x)

OpenStudy (anonymous):

I know the period is 2pi and the range is all real numbers already

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

So what are you missing and what do you think they might be? Make your best guess and I'll help you get it right!

OpenStudy (anonymous):

I am missing domain, asymptotes, and zeros. Asymptotes = 2, pi/2? I really don't know those 3 ones

OpenStudy (anonymous):

@OrangeMaster If you don't know an answer, it's better if you don't reply at all.

OpenStudy (anonymous):

Lol okay

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

y = -3tan(x/2). Period: Since tan(x) has a period of pi, tan(x/2) will have a period of pi / (1/2) = 2(pi). Domain: tan(x) is undefined when x is: pi/2 + n(pi) where n is an integer ...-2, -1, 0, 1, 2, .... Therefore tan(x/2) will be undefined when x is: pi + 2n(pi) or (2n+1)pi. Domain of tan(x/2) is: {x|x≠(2n+1)π,n=...,−2,−1,0,1,...}={x|x≠...,−3π,−π,π,3π,...} Range: tan(x) has the range (-infinity, infinity) and so does tan(x/2) Zeros: tan(x) has zeros at x = n(pi) where n is an integer: ...-1, 0, 1, ... Therefore, tan(x/2) has zeros at x = 2n(pi) where n is an integer: ...-1, 0, 1, ... or at x = ..., -4pi, -2pi, 0, 2pi, 4pi, ... Asymptotes: Vertical asymptotes will be when tan(x/2) goes to +/- infinity. We found those point earlier and excluded them from the domain because tan(x/2) will be undefined at those x values. Those same x values will be the location of the vertical asymptotes. At x = (2n+1)pi where n is an integer ot at x = ..., -3pi, -pi, pi, 3pi, ...

OpenStudy (anonymous):

@Nurali the equation is y=-tan(1/2x). so to recap: Domain= -2,-1,0,1, -3pi, -pi, pi, 3pi Zeros= -4pi,, -2pi, 0, 2pi, 4pi Asymptotes= -3pi, =pi, pi, 3i Is that exactly what I would type?

OpenStudy (anonymous):

I mean y=-3tan(1/2x). Is the recap correct @Nurali

OpenStudy (anonymous):

@primeralph

OpenStudy (primeralph):

Domain is all real numbers.

OpenStudy (anonymous):

so the domain and range are both all real numbers for y=-3tan(1/2x)?

OpenStudy (primeralph):

Think of a number for which you can't find the tangent or the inverse tangent.

OpenStudy (anonymous):

ok that makes sense. what about zeros and asymptotes? I've been trying to get help for almost 2 hours, so thank you!

OpenStudy (primeralph):

Solve for x when the equation equals 0.

OpenStudy (anonymous):

0, +-2, +-2pi?

OpenStudy (primeralph):

There's more. So simply, 2n*pi, where n is an integer.

OpenStudy (anonymous):

so 0 +-2npi, n+-2npi?

OpenStudy (primeralph):

Just 2npi.

OpenStudy (anonymous):

so the answer for zeros is just 2npi?

OpenStudy (primeralph):

You have to specify that n is any integer.

OpenStudy (anonymous):

so I would exactly say, zeros: "2npi where n is any integer"

OpenStudy (primeralph):

Yeah, that works. Put a comma before "where". And pi -don't write that in words.

OpenStudy (anonymous):

Ok :) asymptotes?

OpenStudy (primeralph):

Try plotting the graph.

OpenStudy (anonymous):

I did already

OpenStudy (primeralph):

Then you should see the asymptotes.

OpenStudy (anonymous):

I dont? 2pi?

OpenStudy (primeralph):

What's an asymptote?

OpenStudy (anonymous):

a line that approaches a curve but doesn't meet it

OpenStudy (primeralph):

There's a lot of them in that graph.

OpenStudy (anonymous):

I know the graph because I graphed it already. pi/2?

OpenStudy (anonymous):

is it 0±2πn,π±2πn

OpenStudy (anonymous):

and I specify again that n is any integer?

OpenStudy (primeralph):

Again, if you graphed it right, you'd see the correct asymptotes.

OpenStudy (anonymous):

what are they then? I seriously don't know. I thought they were my last answer

OpenStudy (primeralph):

(2n-1)*pi.

OpenStudy (primeralph):

You should really learn this stuff.

OpenStudy (anonymous):

can you explain why that is the correct answer for asymptotes?

OpenStudy (anonymous):

@primeralph

OpenStudy (primeralph):

For the tangent function, asymptotes occur when tan goes to +-infinity as seen in the graph. This occurs at pi, 3pi, 5pi, and so on. To model odd consecutive numbers, you have to have 2n-1.

OpenStudy (primeralph):

Or 2n+1. Either way.

OpenStudy (anonymous):

ohhh but is (2n-1)*pi the right answer? or 2n+1?

OpenStudy (primeralph):

Both are right. They both model odd numbers.

OpenStudy (anonymous):

should I say (2n+-1)*pi?

OpenStudy (primeralph):

You can pick one or both.

OpenStudy (anonymous):

oh so (2n+-1)* pi is an acceptable answer then? I just don't want to screw up later on

OpenStudy (anonymous):

Actually, I'll just stay with (2n-1)* pi

OpenStudy (primeralph):

Okay, good luck.

OpenStudy (anonymous):

thanks. the test is next week, so hopefully you were right on these or I just completely learned the wrong way lol

OpenStudy (primeralph):

You're in charge of how you learn.

OpenStudy (anonymous):

but your answers and explanation were correct right lol?

OpenStudy (primeralph):

How about you check them out? If you got a test, you should be able to know a right answer from a wrong one by now.

OpenStudy (anonymous):

why would you help me by giving me the wrong answers?

OpenStudy (primeralph):

Then why are you asking if they're right?

OpenStudy (anonymous):

I'm tired lol thanks for the help!

OpenStudy (primeralph):

Good luck. Study well.